# Solitary cluster waves in periodic potentials: Formation, propagation, and soliton-mediated particle transport

### v2

@@ 3 Particle clusters in tilted periodic potential
and if the $n$-cluster does not fragment, i.e. if the non-splitting conditions (8) are obeyed.
@@ 4.2 Basic stable cluster size $n_{\rm b}$ for $f>0$
Interestingly, our simulations show that Eq. (17) remains valid for $f>0$, with $\sigma$ replacing $\sigma_{p,q}$. Hence the residual free space of a stabilizable cluster with size $n<n_{\rm max}(\sigma,f)$ is smaller than that of clusters with size $j<n$. Similarly as in Eq. (19),
$n_{\rm b}(\sigma,f)=\operatorname*{argmin}_{n\in\{1,\ldots,n^{\prime}_{\rm max}\}}(\lceil n\sigma\rceil-n\sigma)\,,$ (26)
where $n^{\prime}_{\rm max}\leq n_{\rm max}(\sigma,f)$.
Equations (25) and (26) imply that $n_{\rm b}(\sigma,f)$ can be determined for any $\sigma<1$ and $0<f<f_{\rm c}$ by the following method: First one checks whether an $n_{\rm max}$-cluster, with $n_{\rm max}(\sigma,f)$ from Eq. (25), is stabilizable by using Eq. (8). If it is not stabilizable, $n$ is decreased by one and the stabilizability of this $n$-cluster is checked. The procedure is repeated until the cluster of size $n$ is stabilizable. This $n$ is equal to $n_{\rm b}$, since by decreasing the cluster size, the residual free space increases.
@@ 10 Soliton-mediated particle currents
$d=\frac{n_{\rm b}\!-\!n_{\rm c}\operatorname*{\,mod\,}n_{\rm b}}{\gcd(n_{\rm b},n_{\rm c})}l_{1}+\frac{n_{\rm c}\operatorname*{\,mod\,}n_{\rm b}}{\gcd(n_{\rm b},n_{\rm c})}l_{2}-k\lceil(n_{\rm b}+n_{\rm c})\sigma\rceil\,,$ (65)
@@ Appendix B Velocity of type B solitons
For deriving $d$, we need to specify positions of the soliton. As initial particle configuration we consider one, where the soliton $n_{\rm c}$-cluster attaches to an $n_{\rm b}$-cluster, see the first line in Figs. B.1(a) and (b). In the respective $n_{\rm b}$-cluster, we tag a particle and calculate how it is displaced after successive soliton circulations. Each circulation is finished when the soliton $n_{\rm c}$-cluster attaches to the $n_{\rm b}$-cluster that contains the tagged particle. Summing up all displacements of the tagged particle in the $k$ soliton circulations, with $k$ from Eq. (64), we obtain $d$.
@@
$x=y+n_{\rm c}\sigma+(s+1)\sigma\,.$ (82)
After one soliton circulation, the position of the soliton changes to $y^{\prime}$, and the position and order number of the tagged particle to $x^{\prime}$ and $s^{\prime}$, $x^{\prime}=y^{\prime}+n_{\rm c}\sigma+(s^{\prime}+1)\sigma$. Accordingly, the tagged particle is displaced by
@@
To derive the displacement $\Delta y$ of the soliton, we note that after each time period $\tau_{\rm sol}$, the soliton is displaced by $\lceil n_{\rm b}\sigma\rceil$. From Eq. (55), we know $N_{\rm b}\lceil n_{\rm b}\sigma\rceil+\lceil(n_{\rm b}+n_{\rm c})\sigma\rceil=L$. Hence, after a circulation, the soliton attaches to the $n_{\rm b}$-cluster containing the tagged particle at a position $y^{\prime}=y+l\lceil n_{\rm b}\sigma\rceil-\lceil(n_{\rm b}+n_{\rm c})\sigma\rceil$, where $l$ is a positive integer that we specify below. The subtraction of the term $\lceil(n_{\rm b}+n_{\rm c})\sigma\rceil$ is due to the fact that the displacement $(y^{\prime}-y)$ after the circulation is larger than $L$ and that positions have to be taken modulo $L$. We thus obtain
@@
Equations (84), (62) and (63) imply that only two displacements $\Delta x_{1,2}$ of the tagged particle can occur after a soliton circulation:
@@
$d$ $=k_{1}\Delta x_{1}+k_{2}\Delta x_{2}$ (90)
$=\frac{n_{\rm b}\!-\!n_{\rm c}\operatorname*{\,mod\,}n_{\rm b}}{\gcd(n_{\rm b},n_{\rm c})}l_{1}+\frac{n_{\rm c}\operatorname*{\,mod\,}n_{\rm b}}{\gcd(n_{\rm b},n_{\rm c})}l_{2}-k\lceil(n_{\rm b}+n_{\rm c})\sigma\rceil\,,$

### v3

@@ 3 Particle clusters in tilted periodic potential
and if the $n$-cluster does not fragment, i.e. if the nonsplitting conditions (8) are obeyed.
@@ 4.2 Basic stable cluster size $n_{\rm b}$ for $f>0$
Interestingly, our simulations show that Eq. (17) remains valid for $f>0$, with $\sigma$ replacing $\sigma_{p,q}$. Hence the residual free space of a stabilizable cluster with size $n$, $n\leq n_{\rm max}(\sigma,f)$, is smaller than that of clusters of size smaller than $n$. Similarly as in Eq. (19),
$n_{\rm b}(\sigma,f)=\operatorname*{argmin}_{n\in\{1,\ldots,n^{\prime}_{\rm max}\}}(\lceil n\sigma\rceil-n\sigma)\,.$ (26)
Here, $n^{\prime}_{\rm max}\leq n_{\rm max}(\sigma,f)$ is the size of the largest stabilizable cluster, i.e. which at a translational stable position according to (12) fulfills the non-splitting conditions (8).
Equations (25) and (26) imply that $n_{\rm b}(\sigma,f)$ can be determined for any $\sigma<1$ and $0<f<f_{\rm c}$ by the following method: First one checks whether the $n_{\rm max}$-cluster, with $n_{\rm max}(\sigma,f)$ from Eq. (25), is stabilizable, i.e. whether it satisfies the non-splitting conditions (8). If it is not stabilizable, $n$ is decreased by one and the stabilizability of this $n$-cluster is checked. The procedure is repeated until the cluster of size $n$ is stabilizable. This $n$ is equal to $n_{\rm b}$, since by decreasing the cluster size, the residual free space increases.
@@ 10 Soliton-mediated particle currents
$d$ $=\frac{n_{\rm b}\!-\!n_{\rm c}\operatorname*{\,mod\,}n_{\rm b}}{\gcd(n_{\rm b},n_{\rm c})}l_{1}\lceil n_{\rm b}\sigma\rceil+\frac{n_{\rm c}\operatorname*{\,mod\,}n_{\rm b}}{\gcd(n_{\rm b},n_{\rm c})}l_{2}\lceil n_{\rm b}\sigma\rceil$
$\hskip 10.00002pt{}-k\lceil(n_{\rm b}+n_{\rm c})\sigma\rceil$ (65)
@@ Appendix B Velocity of type B solitons
For deriving $d$, we need to specify positions of the soliton. As initial particle configuration we consider one, where the soliton $n_{\rm c}$-cluster attaches to an $n_{\rm b}$-cluster, see the first line in Fig. B.1. In the respective $n_{\rm b}$-cluster, we tag a particle and calculate how it is displaced after successive soliton circulations. Each circulation is finished when the soliton $n_{\rm c}$-cluster attaches to the $n_{\rm b}$-cluster that contains the tagged particle. Summing up all displacements of the tagged particle in the $k$ soliton circulations, with $k$ from Eq. (64), we obtain $d$.
@@
$x=y+n_{\rm c}\sigma+s\sigma\,.$ (82)
After one soliton circulation, the position of the soliton changes to $y^{\prime}$, and the position and order number of the tagged particle to $x^{\prime}$ and $s^{\prime}$, $x^{\prime}=y^{\prime}+n_{\rm c}\sigma+s^{\prime}\sigma$. Accordingly, the tagged particle is displaced by
@@
To derive the change $\Delta y$ of the soliton position, we note that after each time period $\tau_{\rm sol}$, the soliton is displaced by $\lceil n_{\rm b}\sigma\rceil$. Hence, the distance $\Delta Y$ traveled by the $n_{\rm c}$-cluster in one circulation is an integer multiple of $\lceil n_{\rm b}\sigma\rceil$, $\Delta Y=m\lceil n_{\rm b}\sigma\rceil$, $m\in\mathbb{N}$. Since $m$ must be larger than the number $N_{\rm b}$ of $n_{\rm b}$-clusters, we can write $m\hskip 0.50003ptn_{\rm b}=N_{\rm b}n_{\rm b}+l\hskip 0.50003ptn_{\rm b}$, where $l$ is a positive integer. The change $\Delta y$ of the soliton position is equal to its distance traveled modulo $L$, $\Delta y=\Delta Y\operatorname*{\,mod\,}L=m\lceil n_{\rm b}\sigma\rceil\operatorname*{\,mod\,}L$. Using $N_{\rm b}\lceil n_{\rm b}\sigma\rceil+N_{\rm sol}\lceil(n_{\rm b}+n_{\rm c})\sigma\rceil=L$ from Eq. (55) with $N_{\rm sol}=1$, we thus obtain
@@
Equations (83), (84), (62) and (63) imply that only two displacements $\Delta x_{1,2}$ of the tagged particle can occur after a soliton circulation:
@@
$d$ $=k_{1}\Delta x_{1}+k_{2}\Delta x_{2}$
$=\frac{n_{\rm b}\!-\!n_{\rm c}\operatorname*{\,mod\,}n_{\rm b}}{\gcd(n_{\rm b},n_{\rm c})}l_{1}\lceil n_{\rm b}\sigma\rceil+\frac{n_{\rm c}\operatorname*{\,mod\,}n_{\rm b}}{\gcd(n_{\rm b},n_{\rm c})}l_{2}\lceil n_{\rm b}\sigma\rceil$
$\hskip 10.00002pt{}-k\lceil(n_{\rm b}+n_{\rm c})\sigma\rceil\,,$ (90)